OpenStudy (anonymous):

Is f(x)= x^-2 an even function? And why?

OpenStudy (kirbykirby):

If a function is even, then \(f(-x)=f(x)\). That is: Check is the following true?: \((-x)^{-2}=x^{-2}\) ?

OpenStudy (anonymous):

Thanks What about f(x): 1- ^3radical x?

OpenStudy (kirbykirby):

\(f(x)=1-\sqrt[3]{x}\) ?

OpenStudy (kirbykirby):

regardless of your function though, to check if it's even, you always check if the relation: \(f(-x) = f(x)\) is true.. that is, if you substitute x with "-x", do you get the same function again?

OpenStudy (anonymous):

Ok, its odd

OpenStudy (anonymous):

Thanks

OpenStudy (kirbykirby):

your welcome .. but becareful.. if a funtion is not even, it doesn't mean it's odd

OpenStudy (kirbykirby):

function*

OpenStudy (kirbykirby):

Just as an example, \(f(x)=x^3\) is odd since \(f(-x)=-f(x)\) \((-x)^3=-x^3\) is true But \(f(x)=1-x^3\) is not odd since \(f(-x)=1-(-x)^3=1+x^3\ne-(1-x^3)\)

OpenStudy (anonymous):

Ok yea i know! But how does it look graphed?

OpenStudy (anonymous):

How do input in my calc?

OpenStudy (kirbykirby):

Oh I am not sure. I never had a graphing calculator so I don't know the inputs.